THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 120 — 1 December 2012

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 120^{th} issue of the Magellanic Clouds Newsletter. It contains many very interesting contributions to a wide range of fields of research – enjoy the read! The next issue is planned to be distributed on the 1^{st} of February, 2013. With the best wishes for the New Year,

Editorially Yours, Jacco van Loon

Refereed Journal Papers

The discovery of a rare WO-type Wolf–Rayet star in the Large Magellanic Cloud

Kathryn F. Neugent^{1}, Philip Massey^{1} and Nidia Morrell^{2}

1Lowell Observatory, USA 2Las Campanas Observatory, Carnegie Observatories, Chile

While observing OB stars within the most crowded regions of the Large Magellanic Cloud, we happened upon a new Wolf–Rayet star in Lucke–Hodge 41, the rich OB association that contains S Doradus and numerous other massive stars. At first glance the spectrum resembled that of a WC4 star, but closer examination showed strong O vi λλ3811,34 lines, leading us to classify it as a WO4. This is only the second known WO in the LMC, and the first known WO4 (the other being a WO3). This rarity is to be expected due to these stars’ short lifespans as they represent the most advanced evolutionary stage in a massive star’s life before exploding as SNe. This discovery shows that while the majority of WRs within the LMC have been discovered, there may be a few WRs left to be found.

Accepted for publication in Astronomical Journal Available from arXiv:1210.0062

An infrared study of fullerene planetary nebulæ

D.A. García-Hernández^{1,2}, E. Villaver^{3}, P. García-Lario^{4}, J.A. Acosta-Pulido^{1,2}, A. Manchado^{1,2,5}, L. Stanghellini^{6}, R.A. Shaw^{6} and F. Cataldo^{7,8}

1Instituto de Astrofísica de Canarias (IAC), Spain 2Universidad de La Laguna (ULL), Spain 3Universidad Autónoma de Madrid (UAM), Spain 4Herschel Science Centre. ESAC–ESA, Spain 5CSIC, Spain 6NOAO, USA 7INAF – Osservatorio Astrofisico di Catánia, Italy 8Actinium Chemical Research, Italy

We present a study of sixteen Planetary Nebulæ (PNe) where fullerenes have been detected in their Spitzer Space Telescope spectra. This large sample of objects offers an unique opportunity to test conditions of fullerene formation and survival under different metallicity environments as we are analyzing five sources in our own Galaxy, four in the Large Magellanic Cloud, and seven in the Small Magellanic Cloud. Among the sixteen PNe under study, we present the first detection of C_{60} (possibly also C_{70}) fullerenes in the PN M 1-60 as well as of the unusual ∼ 6.6, 9.8, and 20 μm features (attributed to possible planar C_{24}) in the PN K 3-54. Although selection effects in the original samples of PNe observed with Spitzer may play a potentially significant role in the statistics, we find that the detection rate of fullerenes in C-rich PNe increases with decreasing metallicity (∼ 5% in the Galaxy, ∼ 20% in the LMC, and ∼ 44% in the SMC) and we interpret this as a possible consequence of the limited dust processing occuring in Magellanic Cloud (MC) PNe. cloudy photoionization modeling matches the observed IR fluxes with central stars that display a rather narrow range in effective temperature (∼ 30, 000–45,000 K), suggesting a common evolutionary status of the objects and similar fullerene formation conditions. Furthermore, the data suggest that fullerene PNe likely evolve from lowmass progenitors and are usually of low-excitation. We do not find a metallicity dependence on the estimated fullerene abundances. The observed C_{60} intensity ratios in the Galactic sources confirm our previous finding in the MCs that the fullerene emission is not excited by the UV radiation from the central star. cloudy models also show that lineand wind-blanketed model atmospheres can explain many of the observed [Ne iii]/[Ne ii] ratios by photoionization suggesting that possibly the UV radiation from the central star, and not shocks, are triggering the decomposition of the circumstellar dust grains. With the data at hand, we suggest that the most likely explanation for the formation of fullerenes and graphene precursors in PNe is that these molecular species are built from the photo-chemical processing of a carbonaceous compound with a mixture of aromatic and aliphatic structures similar to that of HAC dust.

Accepted for publication in The Astrophysical Journal Available from arXiv:1210.0216 and from http://rialto.ll.iac.es/folleto/research/preprints/?c=view&pre id=12043

The VLT-FLAMES Tarantula Survey VIII. Multiplicity properties of the O-type star population

H. Sana^{1}, A. de Koter^{1,2}, S. de Mink^{3,4}, P.R. Dunstall^{5}, C.J. Evans^{6}, V. Hénault-Brunet^{7}, J. Maíz Apellániz^{8}, O.H. Ramírez-Agudelo^{1}, W.D. Taylor^{7}, N.R. Walborn^{3}, J.S. Clark^{9}, P.A. Crowther^{10}, A. Herrero^{11,12}, M. Gieles^{13},

N. Langer^{14}, D.J. Lennon^{15,3} and J.S. Vink^{16}

1Astronomical Institute Anton Pannekoek, Amsterdam University, Science Park 904, 1098 XH, Amsterdam, The Netherlands 2Utrecht University, Princetonplein 5, 3584CC, Utrecht, The Netherlands 3Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 4Johns Hopkins University, 11100 Johns Hopkins Rd. Laurel, Baltimore, MD, USA 5Astrophysics Research Centre, School of Mathematics and Physics, Queens University Belfast, Belfast BT7 1NN, UK 6UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 7Scottish Universities Physics Alliance (SUPA), Institute for Astronomy, University of Edinburgh, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 8Instituto de Astrofísica de Andalucía–CSIC, Glorieta de la Astronomía s/n, E-18008 Granada, Spain 9Department of Physics and Astronomy, The Open University, Walton Hall, Milton Keynes, MK7 6AA, UK 10Dept. of Physics & Astronomy, Hounsfield Road, University of Sheffield, UK 11Instituto de Astrofísica de Canarias, C/ Vía Láctea s/n, E-38200 La Laguna, Tenerife, Spain 12Departamento de Astrofísica, Universidad de La Laguna, Avda. Astrofísico Francisco Sánchez s/n, E-38071 La Laguna, Tenerife, Spain 13Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK 14Argelander-Institut für Astronomie, Universität Bonn, Auf dem Hügel 71, 53121 Bonn, Germany 15ESA/STScI, 3700 San Martin Drive, Baltimore, MD 21218, USA 16Armagh Observatory, College Hill, Armagh, BT61 9DG, Northern Ireland, UK

Context: The Tarantula Nebula in the Large Magellanic Cloud is our closest view of a starburst region and is the ideal environment to investigate important questions regarding the formation, evolution and final fate of the most massive stars. Aims: We analyze the multiplicity properties of the massive O-type star population observed through multi-epoch spectroscopy in the framework of the VLT-FLAMES Tarantula Survey. With 360 O-type stars, this is the largest homogeneous sample of massive stars analyzed to date. Methods: We use multi-epoch spectroscopy and variability analysis to identify spectroscopic binaries. We also use a Monte-Carlo method to correct for observational biases. By modeling simultaneously the observed binary fraction, the distributions of the amplitudes of the radial velocity variations and the distribution of the time scales of these variations, we derive the intrinsic current binary fraction and period and mass-ratio distributions. Results: We observe a spectroscopic binary fraction of 0.35±0.03, which corresponds to the fraction of objects displaying statistically significant radial velocity variations with an amplitude of at least 20 km s^{−1}. We compute the intrinsic binary fraction to be 0.51 ± 0.04. We adopt power-laws to describe the intrinsic period and mass-ratio distributions: f(log_{10} P/d)^{π} (with log_{10} P/d in the range 0.15–3.5) and f(q) ∼ q^{κ} with 0.1 ≤ q = M_{2}/M_{1} ≤ 1.0. The power-law indexes that best reproduce the observed quantities are π = −0.45 ± 0.30 and κ = −1.0 ± 0.4. The period distribution that we obtain thus favours shorter period systems compared to an Öpik law (π = 0). The mass ratio distribution is slightly skewed towards low mass ratio systems but remains incompatible with a random sampling of a classical mass function (κ = −2.35). The binary fraction seems mostly uniform across the field of view and independent of the spectral types and luminosity classes. The binary fraction in the outer region of the field of view (r > 7.′8, i.e. ≈ 117 pc) and among the O9.7 I/II objects are however significantly lower than expected from statistical fluctuations.

The observed and intrinsic binary fractions are also lower for the faintest objects in our sample (K_{s} > 15.5 mag), which results from observational effects and the fact that our O star sample is not magnitude-limited but is defined by a spectral-type cutoff. We also conclude that magnitude-limited investigations are biased towards larger binary fractions. Conclusions: Using the multiplicity properties of the O stars in the Tarantula region and simple evolutionary considerations, we estimate that over 50% of the current O star population will exchange mass with its companion within a binary system. This shows that binary interaction is greatly affecting the evolution and fate of massive stars, and must be taken into account to correctly interpret unresolved populations of massive stars.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1209.4638

Physical conditions in the gas phases of the giant H ii region LMC-N 11 unveiled by Herschel – I. Diffuse [C ii] and [O iii] emission in LMC-N 11B

V. Lebouteiller^{1}, D. Cormier^{1}, S.C. Madden^{1}, F. Galliano^{1}, R. Indebetouw^{2}, N. Abel^{3}, M. Sauvage^{1}, S. Hony^{1}, A. Contursi^{4}, A. Poglitsch^{4}, A. Rémy^{1}, E. Sturm^{4} and R. Wu^{1}

1Laboratoire AIM, CEA/DSM-CNRS-Université Paris Diderot DAPNIA/Service d’Astrophysique Bât. 709, CEA-Saclay F-91191 Gif-sur-Yvette Cédex, France 2Department of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903, USA 3Department of Physics, University of Cincinnati, Cincinnati, OH 45221, USA 4Max-Planck-Institute for Extraterrestrial Physics (MPE), Gießenbachstraße 1, 85748 Garching, Germany

Context: The Magellanic Clouds provide a nearby laboratory for metal-poor dwarf galaxies. The low dust abundance enhances the penetration of UV photons into the interstellar medium (ISM), resulting in a relatively larger filling factor of the ionized gas. Furthermore, there is likely a molecular gas reservoir probed by the [C ii] 157 μm line not traced by CO(1–0), the so-called ”dark” gas. The Herschel Space Telescope allows us to observe far-infrared (FIR) cooling lines and to examine the physical conditions in the gas phases of a low-metallicity environment to unprecedented small spatial scales. Aims: Our objective is to interpret the diffuse emission of FIR cooling lines in the H ii region N 11B in the Large Magellanic Cloud. We first study the distribution and properties of the ionized gas. We then constrain the origin of the [C ii] line by comparing to tracers of the low-excitation ionized gas and of photodissociation regions (PDRs). Methods: We present Herschel/PACS maps of N 11B in several tracers, [C ii] 157 μm, [O i] 63 μm and 145 μm, [N ii] 122 μm, [N iii] 57 μm, and [O iii] 88 μm. Optical images in Hα and [O iii] 5007 Å are used as complementary data to investigate the effect of dust extinction. Observations are interpreted with photoionization models to infer the gas conditions and estimate the ionized gas contribution to the [C ii] emission. PDRs are probed through polycyclic aromatic hydrocarbons (PAHs) observed with the Spitzer Space Telescope. Results: [O iii] 88 μm is dominated by extended emission from the high-excitation diffuse ionized gas; it is the brightest FIR line throughout N 11B, ∼ 4 times brighter than [C ii]. We find that about half of the emission from the ionized gas is extinguished by dust. We modeled [O iii] around each O-type star and find that the density of the ISM is ∼ 16 cm^{−3} on large scales. The extent of the [O iii] emission suggests that the medium is rather fragmented, allowing far-UV photons to permeate into the ISM to scales of ∼ 30 pc. Furthermore, by comparing [C ii] with [N ii] 122 μm, we find that 95% of [C ii] arises in PDRs, except toward the stellar cluster for which as much as 15% could arise in the ionized gas. We find a remarkable correlation between [C ii]+[O i] and PAH emission, with [C ii] dominating the cooling in diffuse PDRs and [O i] dominating in the densest PDRs. The combination of [C ii] and [O i] provides a proxy for the total gas cooling. Our results suggest that PAH emission describes better the gas heating in PDRs as compared to the total infrared emission.

Accepted for publication in A&A Available from arXiv:1209.6057

XMM–Newton view of the N 206 superbubble in the Large Magellanic Cloud

Patrick J. Kavanagh^{1}, Manami Sasaki^{1} and Sean D. Points^{2}

1Institut für Astronomie und Astrophysik, Kepler Center for Astro and Particle Physics, Eberhard Karls Universität, 72076 Tübingen, Germany 2Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile

We perform an analysis of the X-ray superbubble in the N 206 H ii region in the Large Magellanic Cloud using current generation facilities to gain a better understanding of the physical processes at work in the superbubble and to improve our knowledge of superbubble evolution. We used XMM–Newton observations of the N 206 region to produce images and extract spectra of the superbubble diffuse emission. Morphological comparisons with Hα images from the Magellanic Cloud Emission Line Survey were performed, and spectral analysis of the diffuse X-ray emission was carried out. We derived the physical properties of the hot gas in the superbubble based on the results of the spectral analysis. We also determined the total energy stored in the superbubble and compared this to the expected energy input from the stellar population to assess the superbubble growth rate discrepancy for N 206. We find that the brightest region of diffuse X-ray emission is confined by a Hα shell, consistent with the superbubble model. In addition, faint emission extending beyond the Hα shell was found, which we attribute to a blowout region. The spectral analysis of both emission regions points to a hot shocked gas as the likely origin of the emission. We determine the total energy stored in the bubble and the expected energy input by the stellar population. However, due to limited data on the stellar population, the input energy is poorly constrained and, consequently, no definitive indication of a growth rate discrepancy is seen. Using the high-sensitivity X-ray data from XMM–Newton and optical data from the Magellanic Cloud Emission Line Survey has allowed us to better understand the physical properties of the N 206 superbubble and address some key questions of superbubble evolution.

Accepted for publication in A&A Available from arXiv:1209.5247

Tagging the chemical evolution history of the Large Magellanic Cloud disk

Emilio Lapenna^{1}, Alessio Mucciarelli^{1}, Livia Origlia^{2} and Francesco R. Ferraro^{1}

1Department of Astronomy, University of Bologna, Via Ranzani, 1-40127, Bologna, Italy 2INAF–OABo, Via Ranzani, 1-40127, Bologna, Italy

We have used high-resolution spectra obtained with the multifiber facility FLAMES at the Very Large Telescope of the European Southern Observatory to derive kinematic properties and chemical abundances of Fe, O, Mg and Si for 89 stars in the disk of the Large Magellanic Cloud (LMC). The derived metallicity and [α/Fe], obtained as the average of O, Mg and Si abundances, allow us to draw a preliminary scheme of the star formation history occurred in this region of the LMC. The derived metallicity distribution shows two main components: one component (comprising ∼ 84% of the sample) is peaked at [Fe/H] = −0.48 dex and it shows an [α/Fe] ratio slightly under solar ([α/Fe] ∼ −0.1 dex). This population was probably originated by the main star formation event occurred 3–4 Gyr ago (possibly triggered by tidal capture of the Small Magellanic Cloud). The other component (comprising ∼ 16% of the sample) is peaked at [Fe/H] ∼ −1 dex and it shows an [α/Fe] ∼ 0.2 dex. This population was probably generated during the long quiescent epoch of star formation in between the first episode and the most recent bursts. Indeed, in our sample we do not find stars with chemical properties similar to the old LMC globular clusters nor to the iron-rich and α-poor stars recently found in the LMC globular cluster NGC 1718 and predicted to be also in the LMC field, thus suggesting that both these components are small (< 1%) in the LMC disk population.

Accepted for publication in ApJ Available from arXiv:1210.0560

The VLT-FLAMES Tarantula Survey. IX. The interstellar medium seen through Diffuse Interstellar Bands and neutral sodium

Jacco Th. van Loon^{1}, Mandy Bailey^{1}, Benjamin L. Tatton^{1}, Jesús Maíz Apellániz^{2}, Paul A. Crowther^{3}, Alex de Koter^{4}, Christopher J. Evans^{5}, Vincent Hénault-Brunet^{6}, Ian D. Howarth^{7}, Philipp Richter^{8}, Hugues Sana^{4}, Sergio

Simón-Díaz^{9,10}, William Taylor^{5} and Nolan R. Walborn^{11}

1Astrophysics Group, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 2Instituto de Astrofísica de Andalucía–CSIC, Glorieta de la Astronomía s/n, 18008 Granada, Spain 3Department of Physics and Astronomy, University of Sheffield, Hicks Building, Hounsfield Road, Sheffield S3 7RH, UK 4Astronomical Institute Anton Pannekoek, University of Amsterdam, P.O. Box 94249, 1090 GE Amsterdam, The Netherlands 5UK Astronomy Technology Centre, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 6Scottish Universities Physics Alliance (SUPA), Institute for Astronomy, University of Edinburgh, Royal Observatory Edinburgh, Blackford Hill, Edinburgh, EH9 3HJ, UK 7Department of Physics and Astronomy, University College London, Gower Street, London WC1E 6BT, UK 8Institut für Physik und Astronomie, Universität Potsdam, Haus 28, Karl-Liebknecht-Straße 24/25, D-14476 Potsdam, Germany 9Instituto de Astrofísica de Canarias, E-38200 La Laguna, Tenerife, Spain 10Departamento de Astrofísica, Universidad de La Laguna, E-38205 La Laguna, Tenerife, Spain 11Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD, 21218, USA

Context: The Tarantula Nebula (a.k.a. 30 Dor) is a spectacular star-forming region in the Large Magellanic Cloud (LMC), seen through gas in the Galactic Disc and Halo. Diffuse Interstellar Bands (DIBs) offer a unique probe of the diffuse, cool–warm gas in these regions. Aims: The aim is to use DIBs as diagnostics of the local interstellar conditions, whilst at the same time deriving properties of the yet-unknown carriers of these enigmatic spectral features. Methods: Spectra of over 800 early-type stars from the Very Large Telescope Flames Tarantula Survey (VFTS) were analysed. Maps were created, separately, for the Galactic and LMC absorption in the DIBs at 4428 and 6614 Å and – in a smaller region near the central cluster R 136 – neutral sodium (the Na i D doublet); we also measured the DIBs at 5780 and 5797 Å. Results: The maps show strong 4428 and 6614 Å DIBs in the quiescent cloud complex to the south of 30 Dor but weak absorption in the harsher environments to the north (bubbles) and near the OB associations. The Na maps show at least five kinematic components in the LMC and a shell-like structure surrounding R 136, and small-scale structure in the Milky Way. The strengths of the 4428, 5780, 5797 and 6614 Å DIBs are correlated, also with Na absorption and visual extinction. The strong 4428 Å DIB is present already at low Na column density but the 6614, 5780 and 5797 Å DIBs start to be detectable at subsequently larger Na column densities. Conclusions: The carriers of the 4428, 6614, 5780 and 5797 Å DIBs are increasingly prone to removal from irradiated gas. The relative strength of the 5780 and 5797 Å DIBs clearly confirm the Tarantula Nebula as well as Galactic high-latitude gas to represent a harsh radiation environment. The resilience of the 4428 Å DIB suggests its carrier is large, compact and neutral. Structure is detected in the distribution of cool–warm gas on scales between one and > 100 pc in the LMC and as little as 0.01 pc in the Sun’s vicinity. Stellar winds from the central cluster R 136 have created an expanding shell; some infalling gas is also detected, reminiscent of a galactic “fountain”.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1210.4447

Theoretical fit of Cepheid light an radial velocity curves in the Large Magellanic Cloud cluster NGC 1866

Marcella Marconi^{1}, Roberto Molinaro^{1}, Vincenzo Ripepi^{1}, Ilaria Musella^{1} and Enzo Brocato^{2}

1INA–Osservatorio Astronomico di Capodimonte, Napoli, Italy 2INAF–Osservatorio Astronomico di Monteporzio, Roma, Italy

We present a theoretical investigation of multifilter (U,B,V, I and K) light and radial velocity curves of five Classical Cepheids in NGC 1866, a young massive cluster of the Large Magellanic Cloud. The best fit models accounting for the luminosity and radial velocity variations of the five selected variables, four pulsating in the fundamental mode and one in the first overtone, provide direct estimates of their intrinsic stellar parameters and individual distances. The resulting stellar properties indicate a slightly brighter Mass–Luminosity relation than the canonical one, possibly due to mild over-shooting and/or mass loss. As for the inferred distances, the individual values are consistent within the uncertainties. Moreover, their weighted mean value corresponds to a distance modulus of 18.56 ± 0.03 (stat) ±0.1 (syst) mag, in agreement with several independent results in the literature.

Accepted for publication in MNRAS Available from arXiv:1210.4343

Physical parameters and evolutionary route for the LMC interacting binary OGLE 05155332−6925581

Hernán Garrido^{1,2}, Ronald Mennickent^{2}, Gojko Djurašević^{3,4}, Zbigniew Kołaczkowski^{5}, Ewa Niemzcura^{5} and Nicky Mennekens^{6}

1European Organisation for Astronomical Research in the Southern Hemisphere, Chile 2Universidad de Concepción, Chile 3Astronomical Observatory, Volgina, Belgrade, Serbia 4Isaac Newton Institute of Chile, Yugoslavia Branch, Belgrade, Serbia 5Instytut Astronomiczny Uniwersytetu Wrocławskiego Ul. Kopernika, Wrocław, Poland 6Astrophysical Institute, Vrije Universiteit Brussel, Brussels, Belgium

We analyze multicolor light curves and high resolution optical spectroscopy of the eclipsing binary and Double Periodic Variable OGLE 05155332−6925581. According to Mennickent et al., this system shows a significant change in the long non-orbital photometric cycle, a loop in the color–magnitude diagram during this cycle and discrete spectral absorption components that were interpreted as evidence of systemic mass loss. We find that the best fit to the multi-band light curves requires a circumprimary optically thick disc with a radius about twice the radius of the more massive star. The spectroscopy indicates a mass ratio of 0.21 ± 0.02 and masses for the hot and cool stars of 9.1 ± 0.5 and 1.9 ± 0.2 M_{⊙}, respectively. A comparison with synthetic binary-star evolutionary models indicates that the system has an age of 4.76 × 10^{7} years, is in the phase of rapid mass transfer, the second one in the life of this binary, in a Case-B massexchange stage. Donor-subtracted Hα profiles show the presence of double emission formed probably in an optically thin circumstellar medium, while the variable He i profile and the Hβ absorption wings are probably formed in the optically thick circumprimary disc. The model that best fit the observations shows the system with a relatively large mass transfer rate of Ṁ = 3.1 × 10^{−6} M_{⊙} yr^{−1}. However, the orbital period remains relatively stable during almost 15 years. This observation suggests that the hot-spot mass-loss model proposed by other authors is not adequate in this case, and that some other mechanism is efficiently removing angular momentum from the binary. Furthermore, our observations suggest that the DPV phenomenon could have an important effect in the balance of mass and angular momentum in the system.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1210.1227

AKARI infrared camera survey of the Large Magellanic Cloud. I. Point source catalog

Daisuke Kato^{1,2,3}, Yoshifusa Ita^{4}, Takashi Onaka^{1}, Tanabe Toshihiko^{5}, Takashi Shimonishi^{1,10}, Itsuki Sakon^{1}, Hidehiro Kaneda^{6}, Akiko Kawamura^{7}, Takehiko Wada^{2}, Fumihiko Usui^{2}, Bon-Chul Koo^{8}, Mikako Matsuura^{9} and

Hidenori Takahashi^{5}

1Department of Astronomy, Graduate School of Science, The University of Tokyo, Japan 2Institute of Space and Astronautical Science, Japan Aerospace Exploration Agency, Japan 3Center for Low Carbon Society Strategy, Japan Science and Technology Agency, Japan 4Astronomical Institute, Tohoku University, Japan 5Institute of Astronomy, School of Science, The University of Tokyo, Japan 6Department of Astrophysics, Nagoya University, Japan 7National Astronomical Observatory of Japan, Japan 8Department of Physics and Astronomy, Seoul National University, South Korea 9Department of Physics and Astronomy, University College London, UK 10Department of Earth and Planetary Sciences, Graduate School of Science, Kobe University, Japan

We present a near- to mid-infrared point source catalog of 5 photometric bands at 3.2, 7, 11, 15 and 24 μm for a 10 deg^{2} area of the Large Magellanic Cloud (LMC) obtained with the Infrared Camera (IRC) onboard the AKARI satellite. To cover the survey area the observations were carried out at 3 separate seasons from 2006 May to June, 2006 October to December, and 2007 March to July. The 10-σ limiting magnitudes of the present survey are 17.9, 13.8, 12.4, 9.9, and 8.6 mag at 3.2, 7, 11, 15 and 24 μm, respectively. The photometric accuracy is estimated to be about 0.1 mag at 3.2 μm and 0.06–0.07 mag in the other bands. The position accuracy is 0.′′3 at 3.2, 7 and 11 μm and 1.′′0 at 15 and 24 μm. The sensitivities at 3.2, 7, and 24 μm are roughly comparable to those of the Spitzer SAGE LMC point source catalog, while the AKARI catalog provides the data at 11 and 15 μm, covering the mid-infrared spectral range contiguously. Two types of catalog are provided: a Catalog and an Archive. The Archive contains all the detected sources, while the Catalog only includes the sources that have a counterpart in the Spitzer SAGE point source catalog. The Archive contains about 650,000, 140,000, 97,000, 43,000, and 52,000 sources at 3.2, 7, 11, 15, and 24 μm, respectively. Based on the catalog, we discuss the luminosity functions at each band, the color–color diagram, and the color–magnitude diagram using the 3.2, 7, and 11 μm band data. Stars without circumstellar envelopes, dusty C-rich and O-rich stars, young stellar objects, and background galaxies are located at distinct regions in the diagrams, suggesting that the present catalog is useful for the classification of objects towards the LMC.

Accepted for publication in The Astronomical Journal Available from arXiv:1210.3465

Chemical evolution of the LMC

Kenji Bekki^{1} and Takuji Tsujimoto^{2}

1UWA, Australia 2NAO, Japan

We adopt a new chemical evolution model for the Large Magellanic Cloud (LMC) and thereby investigate its past star formation and chemical enrichment histories. The delay time distribution of type Ia supernovae recently revealed by type Ia supernova surveys is incorporated self-consistently into the new model. The principle results are summarized as follows. The present gas mass fraction and stellar metallicity as well as the higher [Ba/Fe] in metal-poor stars at [Fe/H] < −1.5 can be more self-consistently explained by models with steeper initial mass functions. The observed higher [Mg/Fe] (> 0.3) at [Fe/H] ∼ −0.6 and higher [Ba/Fe] (> 0.5) at [Fe/H] ∼ −0.3 can be due to significantly enhanced star formation about 2 Gyr ago. The observed overall [Ca/Fe]–[Fe/H] relation and remarkably low [Ca/Fe] (< −0.2) at [Fe/H] > −0.6 are consistent with models with short-delay supernova Ia and with the more efficient loss of Ca possibly caused by an explosion mechanism of type II supernovæ. Although the metallicity distribution functions do not show double peaks in the models with a starburst about 2 Gyr ago, they show characteristic double peaks in the models with double starbursts at ∼ 200 Myr and ∼ 2 Gyr ago. The observed apparent dip of [Fe/H] around ∼ 1.5 Gyr ago in the age–metallicity relation can be reproduced by models in which a large amount (∼ 10^{9} M_{⊙}) of metal-poor ([Fe/H] < −1) gas can be accreted onto the LMC.

Accepted for publication in ApJ Available from arXiv:1210.3968

Early-stage young stellar objects in the Small Magellanic Cloud

J.M. Oliveira^{1}, J.Th. van Loon^{1}, G.C. Sloan^{2}, M. Sewiło^{3}, K.E. Kraemer^{4}, P.R. Wood^{5}, R. Indebetouw^{6,7}, M.D. Filipović^{8}, E.J. Crawford^{8}, G.F. Wong^{8}, J.L. Hora^{9}, M. Meixner^{10}, T.P. Robitaille^{11}, B. Shiao^{10} and J.D. Simon^{12}

1School of Physical & Geographical Sciences, Lennard-Jones Laboratories, Keele University, Staffordshire ST5 5BG, UK 2Department of Astronomy, Cornell University, Ithaca, NY 14853, USA 3Department of Physics & Astronomy, The Johns Hopkins University, 3400 N. Charles Street, Baltimore, MD 21218, USA 4Boston College, Institute for Scientific Research, 140 Commonwealth Avenue, Chestnut Hill, MA 02467, USA 5Research School of Astronomy and Astrophysics, Australian National University, Cotter Road, Weston Creek, ACT 2611, Australia 6Department of Astronomy, University of Virginia, P.O. Box 3818, Charlottesville, VA 22903, USA 7National Radio Astronomical Observatory, Charlottesville, VA 22904, USA 8University of Western Sydney, Locked Bag 1797, Penrith South DC, NSW 1797, Australia 9Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, MS-65 Cambridge, MA 02138-1516, USA 10Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 11Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117 Heidelberg, Germany 12Observatories of the Carnegie Institution of Washington, 813 Santa Barbara St., Pasadena, CA 91101, USA

We present new observations of 34 Young Stellar Object (YSO) candidates in the Small Magellanic Cloud (SMC). The photometric selection required sources to be bright at 24 and 70 μm (to exclude evolved stars and galaxies). The anchor of the analysis is a set of Spitzer-IRS spectra, supplemented by groundbased 3–5 μm spectra, Spitzer IRAC and MIPS photometry, near-IR imaging and photometry, optical spectroscopy and radio data. The sources’ spectral energy distributions (SEDs) and spectral indices are consistent with embedded YSOs; prominent silicate absorption is observed in the spectra of at least ten sources, silicate emission is observed towards four sources. Polycyclic Aromatic Hydrocarbon (PAH) emission is detected towards all but two sources. Based on band ratios (in particular the strength of the 11.3-μm and the weakness of the 8.6-μm bands) PAH emission towards SMC YSOs is dominated by predominantly small neutral grains. Ice absorption is observed towards fourteen sources in the SMC. The comparison of H_{2}O and CO_{2} ice column densities for SMC, Large Magellanic Cloud (LMC) and Galactic samples suggests that there is a significant H_{2}O column density threshold for the detection of CO_{2} ice. This supports the scenario proposed by Oliveira et al. (2011), where the reduced shielding in metal-poor environments depletes the H_{2}O column density in the outer regions of the YSO envelopes. No CO ice is detected towards the SMC sources. Emission due to purerotational 0–0 transitions of molecular hydrogen is detected towards the majority of SMC sources, allowing us to estimate rotational temperatures and H_{2} column densities. All but one source are spectroscopically confirmed as SMC YSOs. Based on the presence of ice absorption, silicate emission or absorption, and PAH emission, the sources are classified and placed in an evolutionary sequence. Of the 33 YSOs identified in the SMC, 30 sources populate different stages of massive stellar evolution. The presence of ice- and/or silicate-absorption features indicates sources in the early embedded stages; as a source evolves, a compact H ii region starts to emerge, and at the later stages the source’s IR spectrum is completely dominated by PAH and fine-structure emission. The remaining three sources are classified as intermediate-mass YSOs with a thick dusty disc and a tenuous envelope still present. We propose one of the SMC sources is a D-type symbiotic system, based on the presence of Raman, H and He emission lines in the optical spectrum, and silicate emission in the IRS-spectrum. This would be the first dust-rich symbiotic system identified in the SMC.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1210.5193

RR Lyræ variables in the Small Magellanic Cloud - II. The extended area: chemical and structural analysis

E. Kapakos^{1} and D. Hatzidimitriou^{1}

1Department of Physics – Section of Astrophysics, Astronomy & Mechanics, University of Athens, 157 84 Zografos, Athens, Greece

We have performed the Fourier decomposition analysis of 8- and 13-year V-band light curves of a carefully selected sample of 454 fundamental-mode RR Lyrae variables (RRab type), detected in a ≃ 14 square degree area of the Small Magellanic Cloud (SMC) and listed in the Optical Gravitational Lensing Experiment, phase III, Catalogue of Variable Stars. The Fourier decomposition parameters were used to derive metal abundances and distance moduli, following the methodology described by Kapakos, Hatzidimitriou & Soszyński. The average metal abundance of the RRab stars on the new scale of Carretta et al. was found to be ⟨[F e/H]_{C09}⟩ = −1.69 ± 0.41 dex (std, with a standard error of 0.02 dex). A tentative metallicity gradient of −0.013 ± 0.007 dex kpc^{−1} was detected, with increasing metal abundance towards the dynamical center of the SMC, but selection effects are also discussed. The distance modulus of the SMC was re-estimated and was found to be ⟨μ⟩ = 19.13 ± 0.19 (std) in a distance scale where the distance modulus of the Large Magellanic Cloud (LMC) is μ_{LMC} = 18.52 ± 0.06 (std). The average 1-σ line-of-sight depth was found to be σ_{int} = 5.3 ± 0.4 kpc (std), while spatial variations of the depth were detected. The SMC was found to be deeper in the north–eastern region, while metal richer and metal poorer objects in the sample seem to belong to different dynamical structures. The former have smaller scale height and may constitute a thick disk, its width being 10.40 ± 0.02 kpc, and a bulge whose size (radius) is estimated to be 2.09 ± 0.81 kpc. The latter seem to belong to a halo structure with a maximum depth along the line of sight extending over 16 kpc in the SMC central region and falling to ∼ 12 kpc in the outer regions.

Published in Monthly Notices of the Royal Astronomical Society Available from arXiv:1210.1954 and from http://onlinelibrary.wiley.com/doi/10.1111/j.1365-2966.2012.21834.x/full

A luminous Be+white dwarf supersoft source in the Wing of the SMC: MAXI J0158−744

K.L. Li^{1}, Albert K. H. Kong^{1,8}, P.A. Charles^{2}, Ting-Ni Lu^{1}, E.S. Bartlett^{2}, M.J. Coe^{2}, V. McBride^{3,4}, A.

Rajoelimanana^{3,4}, A. Udalski^{5}, N. Masetti^{6} and Thomas Franzen^{7}

1Institute of Astronomy & Department of Physics, National Tsing Hua University, Taiwan 2School of Physics & Astronomy, University of Southampton, Highfield, Southampton SO17 1BJ, UK 3South African Astronomical Observatory, P.O. Box 9, Observatory 7935, South Africa 4University of Cape Town, Private Bag X3, Rondebosch 7701, South Africa 5Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland 6INAF – Istituto di Astrofisica Spaziale e Fisica Cosmica di Bologna, via Gobetti 101, 40129 Bologna, Italy 7CSIRO Astronomy and Space Science, Australia 8Kenda Foundation Golden Jade Fellow

We present a multi-wavelength analysis of the very fast X-ray transient MAXI J0158−744, which was detected by MAXI/GSC on 11/11/2011. The subsequent exponential decline of the X-ray flux was followed with Swift observations, all of which revealed spectra with low temperatures (∼ 100 eV) indicating that MAXI J0158−744 is a new Supersoft Source (SSS). The Swift X-ray spectra near maximum show features around 0.8 keV that we interpret as possible absorption from O viii, and emission from O, Fe and Ne lines. We obtained SAAO and ESO optical spectra of the counterpart early in the outburst and several weeks later. The early spectrum is dominated by strong Balmer and He i emission, together with weaker He ii emission. The later spectrum reveals absorption features that indicate a B1/2IIIe spectral type, and all spectral features are at velocities consistent with the SMC. At this distance, it is a luminous SSS (> 10^{37} erg s^{−1}) but whose brief peak luminosity of > 10^{39} erg s^{−1} in the 2–4 keV band makes it the brightest SSS yet seen at ”hard” X-rays. We propose that MAXI J0158−744 is a Be-WD binary, and the first example to possibly enter ULX territory. The brief hard X-ray flash could possibly be a result of the interaction of the ejected nova shell with the B star wind in which the WD is embedded. This makes MAXI J0158−744 only the third Be/WD system in the Magellanic Clouds, but it is by far the most luminous. The properties of MAXI J0158−744 give weight to previous suggestions that SSS in nearby galaxies are associated with early-type stellar systems.

Accepted for publication in ApJ Available from arXiv:1207.5023

Binarity among the Cepheids in the Magellanic Clouds

László Szabados^{1} and Dóa Nehéz^{2}

1Konkoly Observatory, Research Centre for Astronomy and Earth Sciences, Hungarian Academy of Sciences, Budapest, Hungary 2Department of Astronomy, Loránd Eötvös University, Budapest, Hungary

Spectroscopic binarity of the Cepheid variable HV 914 in the Large Magellanic Cloud is pointed out from the published radial velocity observational data. The list of known binaries among Cepheid type variable stars in the Magellanic Clouds is published in tabular form. The census indicates a serious deficiency of Cepheids with known companions as compared with their Galactic counterparts, whose implications are also discussed. A particular amplitude ratio (A_{V_{rad}}/A_{B}) of individual Magellanic Cepheids is studied in order to select promising candidates of spectroscopic binaries worthy of thorough radial velocity studies.

Published in Monthly Notices of the Royal Astronomical Society Available from arXiv:1210.5064

Swift J045106.8−694803; a highly magnetised neutron star in the Large Magellanic Cloud

Helen Klus^{1}, Elizabeth S. Bartlett^{1}, Antony J. Bird^{1}, Malcolm Coe^{1}, Robin Corbet^{2} and Andrzej Udalski^{3}

1The Faculty of Physical and Applied Sciences, University of Southampton, Highfield, Southampton SO17 1BJ, United Kingdom 2University of Maryland Baltimore County, X-ray Astrophysics Laboratory, Mail Code 662, NASA Goddard Space Flight Center, Greenbelt, MD 20771, USA 3Warsaw University Observatory, Aleje Ujazdowskie 4, 00-478 Warsaw, Poland

We report the analysis of a highly magnetised neutron star in the Large Magellanic Cloud (LMC). The high mass X-ray binary pulsar Swift J045106.8−694803 has been observed with Swift X-ray telescope (XRT) in 2008, The Rossi X-ray Timing Explorer (RXTE) in 2011 and the X-ray Multi-Mirror Mission – Newton (XMM–Newton) in 2012. The change in spin period over these four years indicates a spin-up rate of −5.01 ± 0.06 s yr^{−1}, amongst the highest observed for an accreting pulsar. This spin-up rate can be accounted for using Ghosh & Lamb’s (1979) accretion theory assuming it has a magnetic field of (1.2_{−0.7}^{+0.2}) × 10^{14} Gauss. This is over the quantum critical field value. There are very few accreting pulsars with such high surface magnetic fields and this is the first of which to be discovered in the LMC. The large spin-up rate is consistent with Swift Burst Alert Telescope (BAT) observations which show that Swift J045106.8−694803 has had a consistently high X-ray luminosity for at least five years. Optical spectra have been used to classify the optical counterpart of Swift J045106.8−694803 as a B0–1 III–V star and a possible orbital period of 21.631 ± 0.005 days has been found from MACHO optical photometry.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1210.7680

The spatial distribution of dust and stellar emission of the Magellanic Clouds

Ramin A. Skibba^{1,2}, Charles W. Engelbracht^{1,3}, Gonzalo Aniano^{4}, Brian Babler^{5}, Jean-Philippe Bernard^{6,7}, Caroline Bot^{8,9}, Lynn Redding Carlson^{10}, Maud Galametz^{11}, Frederic Galliano^{12}, Karl Gordon^{13}, Sacha Hony^{12}, Frank Israel^{10}, Vianney Lebouteiller^{12}, Aigen Li^{14}, Suzanne Madden^{12}, Margaret Meixner^{13}, Karl Misselt^{1}, Edward

Montiel^{1}, Koryo Okumura^{12}, Pasquale Panuzzo^{12}, Deborah Paradis^{6,7}, Julia Roman-Duval^{13}, Mónica Rubio^{15}, Marc Sauvage^{12}, Jonathan Seale^{13}, Sundar Srinivasan^{16} and Jacco Th. van Loon^{17}

1Steward Observatory, University of Arizona, 933 N. Cherry Avenue, Tucson, AZ 85721, USA 2Center for Astrophysics and Space Sciences, Department of Physics, University of California, 9500 Gilman Dr., La Jolla, San Diego, CA 92093, USA 3Raytheon Company, 1151 East Hermans Road, Tucson, AZ 85756, USA 4Princeton University Observatory, Peyton Hall, Princeton, NJ 08544, USA 5Department of Astronomy, 475 North Charter St., University of Wisconsin, Madison, WI 53706, USA 6Université de Toulouse, UPS-OMP, IRAP, Toulouse, France 7CNRS, IRAP, 9 Av. colonel Roche, BP 44346, F-31028 Toulouse cédex 4, France 8Université de Strasbourg, Observatoire Astronomique de Strasbourg, F-67000 Strasbourg, France 9CNRS, Observatoire Astronomique de Strasbourg, UMR7550, F-67000 Strasbourg, France 10Sterrewacht Leiden, Leiden University, P.O. Box 9513, 2300 RA Leiden, The Netherlands 11Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge CB3 0HA, UK 12AIM, CEA/Saclay, L’Orme des Merisiers, 91191 Gif-sur-Yvette, France 13Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 14314 Physics Building, Department of Physics and Astronomy, University of Missouri, Columbia, MO 65211, USA 15Departamento de Astronomía, Universidad de Chile, Casilla 36-D, Santiago, Chile 16UPMC-CNRS UMR7095, Institut d’Astrophysique de Paris, F-75014 Paris, France 17Lennard-Jones Laboratories, Keele University, ST5 5BG, UK

We study the emission by dust and stars in the Large and Small Magellanic Clouds, a pair of low-metallicity nearby galaxies, as traced by their spatially resolved spectral energy distributions (SEDs). This project combines Herschel Space Observatory PACS and SPIRE far-infrared photometry with other data at infrared and optical wavelengths. We build maps of dust and stellar luminosity and mass of both Magellanic Clouds, and analyze the spatial distribution of dust/stellar luminosity and mass ratios. These ratios vary considerably throughout the galaxies, generally between the range 0.01 ≤ L_{dust}/L_{⋆} ≤ 0.6 and 10^{−4} ≤ M_{dust}/M_{⋆} ≤ 4 × 10^{−3}. We observe that the dust/stellar ratios depend on the interstellar medium (ISM) environment, such as the distance from currently or previously star-forming regions, and on the intensity of the interstellar radiation field (ISRF). In addition, we construct star formation rate (SFR) maps, and find that the SFR is correlated with the dust/stellar luminosity and dust temperature in both galaxies, demonstrating the relation between star formation, dust emission and heating, though these correlations exhibit substantial scatter.

Accepted for publication in Astrophysical Journal Available from arXiv:1210.7812

Diversity of Type Ia supernovæ imprinted in chemical abundances

Takuji Tsujimoto^{1} and Toshikazu Shigeyama^{2}

1National Astronomical Observatory of Japan, Japan 2University of Tokyo, Japan

A time delay of Type Ia supernova (SN Ia) explosions hinders the imprint of their nucleosynthesis on stellar abundances. However, some occasional cases give birth to stars that avoid enrichment of their chemical compositions by massive stars and thereby exhibit an SN Ia-like elemental feature including a very low [Mg/Fe] (∼ −1). We highlight the elemental feature of Fe-group elements for two low-Mg/Fe objects detected in nearby galaxies, and propose the presence of a class of SNe Ia that yield the low abundance ratios of [Cr,Mn,Ni/Fe]. Our novel models of chemical evolution reveal that our proposed class of SNe Ia (slow SNe Ia) is associated with ones exploding on a long timescale after their stellar birth, and give a significant impact on the chemical enrichment in the Large Magellanic Cloud (LMC). In the Galaxy, on the other hand, this effect is unseen due to the overwhelming enrichment by the major class of SNe Ia that explode promptly (prompt SNe Ia) and eject a large amount of Fe-group elements. This nicely explains the different [Cr,Mn,Ni/Fe] features between the two galaxies as well as the puzzling feature seen in the LMC stars exhibiting very low Ca but normal Mg abundances. Furthermore, the corresponding channel of slow SN Ia is exemplified by performing detailed nucleosynthesis calculations in the scheme of SNe Ia resulting from a 0.8+0.6 M_{⊙} white dwarf merger.

Accepted for publication in ApJ Letters Available from arXiv:1210.7829

Milky Way mass models for orbit calculations

Andreas Irrgang^{1}, Benjamin Wilcox^{1,2}, Evan Tucker^{1,3} and Lucas Schiefelbein^{1,4}

1Dr. Karl Remeis-Observatory & ECAP, Astronomical Institute, Friedrich-Alexander University Erlangen-Nüremberg, Sternwartstr. 7, D-96049 Bamberg, Germany 2School of Physics and Astronomy, University of St. Andrews, North Haugh, St. Andrews, KY16 9SS, UK 3Department of Physics, University of Arizona, Tucson, AZ 85721, USA 4Massachusetts Institute of Technology, Cambridge, MA 02139, USA

Context: Studying the trajectories of objects like stars, globular clusters or satellite galaxies in the Milky Way allows to trace the dark matter halo but requires reliable models of its gravitational potential. Aims: Realistic, yet simple and fully analytical models have already been presented in the past. However, improved as well as new observational constraints have become available in the meantime calling for a recalibration of the respective model parameters. Methods: Three widely used model potentials are revisited. By a simultaneous least-squared fit to the observed rotation curve, in-plane proper motion of Sgr A^{⋆}, local mass/surface density and the velocity dispersion in Baade’s window, parameters of the potentials are brought up-to-date. The mass at large radii – and thus in particular that of the dark matter halo – is hereby constrained by imposing that the most extreme halo blue horizontal-branch star known has to be bound to the Milky Way. Results: The Galactic mass models are tuned to yield a very good match to recent observations. The mass of the dark matter halo is – within the limitations of the applied models – estimated in a fully consistent way. As a first application, the trajectory of the hypervelocity star HE 0437−5439 is re-investigated to check its suggested origin in the Large Magellanic Cloud (LMC). Conclusions: Despite their simplicity, the presented Milky Way mass models are well able to reproduce all observational constraints. Their analytical and simple form makes them ideally suited for fast and accurate orbit calculations. The LMC cannot be ruled out as HE 0437−5439’s birthplace.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1211.4353

Variability of the pulsed radio emission from the Large Magellanic Cloud pulsar PSR J0529−6652

Fronefield Crawford^{1}, Dean Altemose^{1}, Hannan Li^{1} and Duncan R. Lorimer^{2}

1Department of Physics and Astronomy, Franklin & Marshall College, P.O. Box 3003, Lancaster, PA 17604, USA 2Department of Physics, West Virginia University, Morgantown, WV 26506, USA

We have studied the variability of PSR J0529−6652, a radio pulsar in the LMC, using observations conducted at 1390 MHz with the Parkes 64-m telescope. PSR J0529−6652 is detectable as a single pulse emitter, with amplitudes that classify the pulses as giant pulses. This makes PS J0529−6652 the second known giant pulse emitter in the LMC, after PSR B0540−69. The fraction of the emitted pulses detectable from PSR J0529−6652 at this frequency is roughly two orders of magnitude greater than it is for either PSR B0540−69 or the Crab pulsar (if the latter were located in the LMC). We have measured a pulse nulling fraction of 83.3 ± 1.5% and an intrinsic modulation index of 4.07 ± 0.29 for PSR J0529−6652. The modulation index is significantly larger than values previously measured for typical radio pulsars but is comparable to values reported for members of several other neutron star classes. The large modulation index, giant pulses, and large nulling fraction suggest that this pulsar is phenomenologically more similar to these other, more variable sources, despite having spin and physical characteristics that are typical of the unrecycled radio pulsar population. The large modulation index also does not appear to be consistent with the small value predicted for this pulsar by a model of polar cap emission outlined by Gil & Sendyk (2000). This conclusion depends to some extent on the assumption that PSR J0529−6652 is exhibiting core emission, as suggested by its simple profile morphology, narrow profile width, and previously measured profile polarization characteristics.

Accepted for publication in Astrophysical Journal Available from arXiv:1211.4531

Giant stellar arcs in the Large Magellanic Cloud: a possible link with past activity of the Milky Way nucleus

Yuri N. Efremov^{1}

1Sternberg Astronomical Institute, MSU, Russia

The origin of the giant stellar arcs in the LMC remains a controversial issue, discussed since 1966. No other stellar arc is so perfect a segment of a circle, moreover, there is another similar arc nearby. Many hypotheses were advanced to explain these arcs, and all but one of these were disproved. It was proposed in 2004 that origin of these arcs was due to the bow shock from the jet, which is intermittently fired by the Milky Way nucleus and during the last episode of its activity the jet was pointed to the LMC. Quite recently evidence for such a jet has really appeared. We suppose it was once energetic enough to trigger star formation in the LMC, and if the jet opening angle was about 2^{∘}, it could push out H i gas from the region of about 2 kpc in size, forming a cavity LMC 4, but also squeezed two dense clouds, which occurred in the same area, causing the formation of stars along their surfaces facing the core of the MW. In result, spherical segments of the stellar shells might arise, visible now as the arcs of Quadrant and Sextant, the apices of which point to the center of the MW. This orientation of both arcs can be the key to unlocking their origin. Here we give data which confirm the above hypothesis, amongst which are radial velocities of stars inside and outside the larger one of the LMC arcs. The probability is low that a jet from an AGN points to a nearby galaxy and triggers star formation there, but a few other examples are now known or suspected.

Accepted for publication in MNRAS Letters Available from arXiv:1211.0684

A VLT/FLAMES study of the peculiar intermediate-age Large Magellanic Cloud star cluster NGC 1846 – I. Kinematics

Dougal Mackey^{1}, Gary Da Costa^{1}, Annette Ferguson^{2} and David Yong^{1}

1RSAA, Australian National University, Australia 2IfA, University of Edinburgh, Scotland

In this paper we present high resolution VLT/FLAMES observations of red giant stars in the massive intermediate-age Large Magellanic Cloud star cluster NGC 1846, which, on the basis of its extended main-sequence turn-off (EMSTO), possesses an internal age spread of ≈ 300 Myr. We describe in detail our target selection and data reduction procedures, and construct a sample of 21 stars possessing radial velocities indicating their membership of NGC 1846 at high confidence. We consider high-resolution spectra of the planetary nebula Mo 17, and conclude that this object is also a member of the cluster. Our measured radial velocities allow us to conduct a detailed investigation of the internal kinematics of NGC 1846, the first time this has been done for an EMSTO system. The key result of this work is that the cluster exhibits a significant degree of systemic rotation, of a magnitude comparable to the mean velocity dispersion. Using an extensive suite of Monte Carlo models we demonstrate that, despite our relatively small sample size and the substantial fraction of unresolved binary stars in the cluster, the rotation signal we detect is very likely to be genuine. Our observations are in qualitative agreement with the predictions of simulations modeling the formation of multiple populations of stars in globular clusters, where a dynamically cold, rapidly rotating second generation is a common feature. NGC 1846 is less than one relaxation time old, so any dynamical signatures encoded during its formation ought to remain present.

Accepted for publication in ApJ Available from arXiv:1211.3146

Multiwavelength study of the newly confirmed supernova remnant MCSNR J0527−7104 in the Large Magellanic Cloud

Patrick J. Kavanagh^{1}, Manami Sasaki^{1}, Sean D. Points^{2}, Miroslav D. Filipović^{3}, Pierre Maggi^{4}, Luke M. Bozzetto^{3}, Evan J. Crawford^{3}, Frank Haberl^{4} and Wolfgang Pietsch^{4}

1Institut für Astronomie und Astrophysik, Kepler Center for Astro and Particle Physics, Eberhard Karls Universität, Sand 1, 72076 Tübingen, Germany 2Cerro Tololo Inter-American Observatory, Casilla 603, La Serena, Chile 3University of Western Sydney, Locked Bag 1797, Penrith NSW 2751, Australia 4Max-Planck-Institut für extraterrestrische Physik, Gießenbachstraße, 85748 Garching, Germany

The Large Magellan ic Cloud (LMC) hosts a rich and varied population of supernova remnants (SNRs). Optical, X-ray, and radio observations are required to identify these SNRs, as well as to ascertain the various processes responsible for the large array of physical characteristics observed. In this paper we attempted to confirm the candidate SNR [HP99] 1234, identified in X-rays with ROSAT, as a true SNR by supplementing these X-ray data with optical and radio observations. Optical data from the Magellanic Cloud Emission Line Survey (MCELS) and new radio data from the Molonglo Observatory Synthesis Telescope (MOST), in addition to the ROSAT X-ray data, were used to perform a multiwavelength morphological analysis of this candidate SNR. An approximately ellipsoidal shell of enhanced [S ii] emission, typical of an SNR ([S ii]/Hα > 0.4), was detected in the optical. This enhancement is positionally coincident with faint radio emission at λ = 36 cm. Using the available data we estimated the size of the remnant to be ∼ 5.′1×4.′0 (∼ 75 pc × 59 pc). However, the measurement along the major-axis was somewhat uncertain due to a lack of optical and radio emission at its extremities and the poor resolution of the X-ray data. Assuming this SNR is in the Sedov phase and adopting the ambient mass density of 1.2 × 10^{−25} g cm^{−3} measured in a nearby H ii region, an age estimate of ∼ 25 kyr was calculated for a canonical initial explosion energy of 10^{51} erg. However, this age estimate should be treated cautiously due to uncertainties on the adopted parameters. Analysis of the local stellar population suggested a type Ia event as a precursor to this SNR, however, a core-collapse mechanism could not be ruled out due to the possibility of the progenitor being a runaway massive star. With the detection of X-ray, radio and significant optical line emission with enhanced [S ii], this object was confirmed as an SNR to which we assign the identifier MCSNR J0527−7104.

Accepted for publication in Astronomy & Astrophysics Available from arXiv:1211.4746

The quasi-resonant variability of the massive LMC system BI 108

Zbigniew Kołaczkowski^{1}, Ronald Mennickent^{2}, Thomas Rivinius^{3} and Grzegorz Pietrzyński^{1,4}

1Instytut Astronomiczny Uniwersytetu Wrocławskiego, Kopernika 11, 51-622 Wrocław, Poland 2Universidad de Concepción, Departamento de Astronomía, Casilla 160-C, Concepción, Chile 3ESO, European Organization for Astronomical Research in the Southern Hemisphere, Chile 4Warsaw University Observatory, Al. Ujazdowskie 4, 00-478 Warszawa, Poland

The early B supergiant LMC star BI 108 is photometrically variable with a unique light curve; two strong periods are present in an almost precise 3:2 resonance. We collected spectroscopic data at VLT/UVES, sampling the supercycle of 10.733 days in ten epochs. We find spectral signatures for a SB2 system consisting of two massive B1 supergiants orbiting at the orbital period of 5.366 days. The shorter periodicity resembles the light curve of an eclipsing binary with periodicity 3.578 days that is not detected in the data. We discuss possible causes for the short periodicity and conclude that the quadruple system is the more plausible hypothesis.

Accepted for publication in Monthly Notices of the Royal Astronomical Society Available from arXiv:1211.4172

GASS High Velocity Clouds in the region of the Magellanic Leading Arm

Bi-Qing For^{1}, Lister Staveley-Smith^{1} and Naomi McClure-Griffiths^{2}

1International Centre for Radio Astronomy Research, University of Western Australia, Australia 2CSIRO Astronomy and Space Science, Australia

We present a catalog of high-velocity clouds in the region of the Magellanic Leading Arm. The catalog is based on neutral hydrogen (H i) from the Parkes Galactic All-Sky Survey (GASS). Excellent spectral resolution allows clouds with narrow-line components to be resolved. The total number of detected clouds is 419. We describe the method of cataloging and present the basic parameters of the clouds.We discuss the general distribution of the high-velocity clouds and classify the clouds based on their morphological type. The presence of a significant number of head– tail clouds and their distribution in the region is discussed in the context of Magellanic System simulations. We suggest that ram-pressure stripping is a more important factor than tidal forces for the morphology and formation of the Magellanic Leading Arm and that different environmental conditions might explain the morphological difference between the Magellanic Leading Arm and Magellanic Stream. We also discuss a newly identified population of clouds that forms the LA IV and a new diffuse bridge-like feature connecting the Leading Arm II and III complexes.

Submitted to ApJ Available from arXiv:1208.5583

Conference Papers

The supershell–molecular cloud connection in the Milky Way and beyond

J.R. Dawson^{1}, N.M. McClure-Griffiths^{2}, Y. Fukui^{3}, J. Dickey^{1}, T. Wong^{4}, A. Hughes^{5} and A. Kawamura^{6}

1School of Mathematics & Physics, University of Tasmania, Sandy Bay, TAS 7005, Australia 2Australia Telescope National Facility, CASS, Marsfield NSW 2122, Australia 3Dept. of Astrophysics, Nagoya University, Chikusa-ku, Nagoya, Japan 4Astronomy Department, University of Illinois, Urbana, IL 61801, USA 5Max-Planck-Institut für Astronomie, Königstuhl 17, D-69117, Heidelberg, Germany 6National Astronomical Observatory of Japan, Tokyo 181-8588, Japan

The role of large-scale stellar feedback in the formation of molecular clouds has been investigated observationally by examining the relationship between H i and ^{12}CO(J = 1–0) in supershells. Detailed parsec-resolution case studies of two Milky Way supershells demonstrate an enhanced level of molecularisation over both objects, and hence provide the first quantitative observational evidence of increased molecular cloud production in volumes of space affected by supershell activity. Recent results on supergiant shells in the LMC suggest that while they do indeed help to organise the ISM into over-dense structures, their global contribution to molecular cloud formation is of the order of only ∼ 10%.

Oral contribution, published in IAUS 292, ”Molecular Gas, Dust, and Star Formation in Galaxies” Available from arXiv:1210.0169

The slow X-ray pulsar SXP 1062 and associated supernova remnant in the Wing of the Small Magellanic Cloud

L.M. Oskinova, L. M.^{1}, M.A. Guerrero, V. Hénault-Brunet, W. Sun, Y.-H. Chu, C.J. Evans, J.S. Gallagher, III, R.A. Gruendl and J. Reyes-Iturbide

1Institute for Physics and Astronomy, University of Potsdam, Germany

SXP 1062 is an exceptional case of a young neutron star in a wind-fed high-mass X-ray binary associated with a supernova remnant. A unique combination of measured spin period, its derivative, luminosity and young age makes this source a key probe for the physics of accretion and neutron star evolution. Theoretical models proposed to explain the properties of SXP 1062 shall be tested with new data.

Poster contribution, published in IAUS 291 ”Neutron Stars and Pulsars: Challenges and Opportunities after 80 years”, ed. J. van Leeuwen Available from arXiv:1210.5066

Review Paper

Circumstellar masers in the Magellanic Clouds

Jacco Th. van Loon^{1}

1Lennard-Jones Laboratories, Keele University, ST5 5BG, UK

The nearby dwarf irregular galaxies the Large and Small Magellanic Clouds have metallicities of about half and a fifth solar, respectively, which offers the unique opportunity to study astrophysical processes as a function of metallicity. Masers in the outflows from evolved stars allow to measure the wind speed, vital to derive mass-loss rates and test wind driving mechanisms. The metallicity dependence of the wind speed in particular allows us to make inferences about dust formation and mass loss in the early Universe. I will review past surveys for circumstellar OH, water, and SiO masers in the Magellanic Clouds (and provide a literature review of interstellar masers). I will then discuss the way these measurements have influenced our understanding of mass loss, and end with outlining the prospects for future surveys for OH masers in the Magellanic Clouds.

Published in ”Science with Parkes @ 50 Years Young”, Ed. Robert Braun (2012) Available from arXiv:1210.0983 and from http://www.atnf.csiro.au/research/conferences/Parkes50th/ProcPapers/vanLoon.pdf